The members of a College staff club were asked to indicate the brand of beer they drink. 5 members drank stars and Trophy beers, 7 members drank star and club beers, 6 members drank starbeer only, 11 drank Trophy beer only while 5 drank club only. Three members drank all the three brands. assuming that each member drank at least one brand of beer, Find
i, the number of members that drank star beer
ii, the number of members that drank Trophy beer
iii, the number of members that drank club beer
iv, the total number of members in the club

Found 2 solutions by math_tutor2020, MathTherapy:
Answer by math_tutor2020(3835)


Here's the venn diagram with 8 distinct regions

The region labels are:
S = people who drank star beer only
T = people who drank trophy beer only
C = people who drank club beer only
ST = people who drank star and trophy, but not club
TC = people who drank trophy and club, but not star
SC = people who drank star and club, but not trophy
STC = people who drank all three brands
N = people who drank none of the three brands mentioned

Given Facts:
NumberFact
15 members drank stars and Trophy beers
27 members drank star and club beers
36 members drank starbeer only
411 drank Trophy beer only
55 drank club only
63 members drank all the three brands
7Each member drank at least one brand of beer

Fact 6 indicates the value "3" goes in region STC.
Facts 1 and 6 will lead to 5-3 = 2 members who drank stars and trophy, but not club beer. Write "2" in region ST
Facts 2 and 6 will lead to 7-3 = 4 members who drank star and club, but not trophy beer. Write "4" in region SC

Fact 3 indicates we'll write "6" in region S.
Fact 4 indicates we'll write "11" in region T.
Fact 5 indicates we'll write "5" in region C.
Fact 7 means we have "0" in region N, because each person drank at least one type.

This is what we have so far

Your teacher doesn't give enough info to determine what value goes in region TC.
I would ask your teacher for clarification about this.

------------------------------------------------------


i, the number of members that drank star beer

We add up the values inside the circle labeled "Star"
6+2+4+3 = 15

Answer: 15 people drank star beer


------------------------------------------------------

ii, the number of members that drank Trophy beer


Like I mentioned earlier, we don't know what value goes in region TC.
We unfortunately do not have enough information to determine how many people drank trophy beer.
Similar situations apply to the remaining questions as well.

Answer by MathTherapy(10809)

The members of a College staff club were asked to indicate the brand of beer they drink. 5 members drank stars and Trophy beers, 7 members drank star and club beers, 6 members drank starbeer only, 11 drank Trophy beer only while 5 drank club only. Three members drank all the three brands. assuming that each member drank at least one brand of beer, Find
i, the number of members that drank star beer
ii, the number of members that drank Trophy beer
iii, the number of members that drank club beer
iv, the total number of members in the club
.
The answers to your questions, i) - iv) are derived from the above Venn Diagram, and are based on the following:
         1) Zero (0) members drink Trophy and Club beers, only.
         2) All members of the club are beer drinkers.

So,   i, the number of members that drank star beer: 6 + 2 + 3 + 4 = 15
     ii, the number of members that drank Trophy beer: 11 + 2 + 3 + 0 = 16
    iii, the number of members that drank club beer: 5 + 4 + 3 + 0 = 12
     iv, the total number of members in the club: 11 + 2 + 3 + 0 + 6 + 4 + 5 = 31